# Large Language Models as Reliable Knowledge Bases?

### v1

@@ 3.2 LLM Selection
We evaluate a wide range of popular LLMs (26 in total) in their ability to function as KBs and investigate the impact of various factors, such as the number of parameters and fine-tuning. Specifically, we consider the following models: gpt-3.5-turbo, flan-t5 (0.08B, 0.25B, 0.78B, 3B, 11B), llama1 (7B, 13B, 65B), llama2 (7B, 13B, 70B), llama2-chat (7B, 13B, 70B), llama3 (8B, 70B), llama3instruct (8B, 70B), mistral (7B), mistral-instruct (7B), gemma (2B, 7B), gemma-instruct (2B, 7B), and Phi2 (3B). Detailed descriptions of these models are provided in Table 4 in Appendix B. We classify LLMs into three categories based on their parameter sizes: small (0.08B–3B), medium (7B–13B), and large (65B–70B). We use the term ‘fine-tuned LLMs’ to refer to the LLMs that have been fine-tuned through instruction-tuning or reinforcement learning from human feedback. We use the term ‘base LLMs’ to refer to LLMs without fine-tuning.
3.3 Evaluation on a Single Response
Uninformative
We identify three types of uninformative responses from LLMs: ‘repetition’, ‘none’, and ‘unsure’. ‘Repetition’ refers to responses that repeatedly echo a specific string. We detect this using regular expressions and word frequency analysis. ‘None’ denotes responses lacking relevant information, such as an empty string and repetition of the question. ‘Unsure’ indicates responses where the model explicitly states it is unable to answer or lacks the required knowledge. We label a response as ‘unsure’ if it includes expressions such as ‘I am not sure’, ‘I cannot provide’, and ‘I am just an AI’.
Correct
We determine whether a response is correct based on exact match. A response is considered correct if the exact match score is 1. In cases where exact match is 0, we compare the model’s prediction against the ground truth using gpt-4o. Previous work Sun et al. (2023) has shown a 98% agreement rate between ChatGPT and human judgments in comparing model responses with ground truth. We follow their prompt as outlined in Table 7 in Appendix C.
Consistency Score
To compute $Cons(q,r)$, we set $X_{\text{MCQs}}$ (total number of MCQs) to 20, and each MCQ includes question $q$ and 5 options (the original response $r$, 3 random distractor options, and an ‘unsure’ option).
3.4 Prompts and Hyper-parameters
All LLMs were evaluated using greedy decoding (temperature 0 for gpt-3.5-turbo) with a maximum of 100 new tokens. Our prompts are provided in Appendix C.
Figure 4: Comparing llama3instruct-8b against llama2chat-13b, and phi2-3b against gemma-2b using CR, WR, and NCR metrics (zero-shot setting). Even though these models are comparable according to CR, they are quite different according to WR and NCR metrics. We observe similar trends in the four-shot and with two unsure shots settings.
4 Results
We present detailed results for all LLMs are in Table 9 (factuality), Table 10 (consistency), and Table 11 (reliability) in Appendix D. LLM rankings based on different metrics are shown in Figure 12 and Figure 13, also in Appendix D.
gpt-3.5-turbo is overall the most reliable LLM.
Table 1 presents results for the two best performing LLMs under different prompt settings. As can be seen, gpt-3.5-turbo is most reliable across the board. Although it is not consistently wrong when asked about facts it does not know (its IUR score exceeds 95%), it is not consistently correct when asked about facts it has seen before (NCCR is only 32%).
Figure 5: LLM ranking based on NCCR and IUR metrics in zero-shot setting. See full results in Figure 12 (Appendix D) for other shot settings.
flan-t5-0.78B is the most reliable LLM for unseen knowledge and most unreliable with seen knowledge.
Figure 5 shows the LLMs ranked by NCCR and IUR in a zero-shot setting. As can be seen, while flan-t5-0.78B ranks low for NCCR, it maintains the top position for IUR. A similar trend is observed with the flan-t5-3B and gemma-instruct (2B, 7B) models. Conversely, models in the llama3 family show the opposite trend: they rank high on seen knowledge but low on unseen knowledge. For instance, llama3-70b, despite ranking second in terms of NCCR, falls to the fourth lowest position for IUR.
Net Correct Rate (NCR) reveals factuality gaps in LLMs.
As illustrated in Figure 4, CR, the standard metric for assessing factuality, fails to fully capture nuanced differences. Despite similar CR values of approximately 40%, llama3instruct-8B and llama2chat-13B behave differently when it comes to wrong responses (on seen data); the former model has a WR of 15% higher than the latter, and as a result its NCR is substantially lower. In the case of phi2-3B, Gunasekar et al. (2023) claim that with "textbook quality" data, smaller LLMs can achieve satisfactory performance with less training data. Focusing solely on CR, this claim is true as phi2-3B’s CR is comparable to that of gemma-2B, which was trained with twice the token count. However, phi2-3B exhibits significantly higher WR (and lower NCR) compared to gemma-2B which underscores the challenge of maintaining low error rates even with superior data quality.
A less factual LLM can be more reliable.
Figure 6 illustrates this finding. The llama3instruct-8b model has a slightly worse NCR compared to llama2chat-7B, indicating that the former model is marginally less factual on seen knowledge. However, llama3instruct-8b exhibits a higher NCCR due to its better consistency score on correct responses. This suggests that llama3instruct-8b is more reliable than llama2chat-7B on seen knowledge. Additionally, both LLMs show negative NCR and positive NCCR in the zero-shot setting. This indicates that although they produce more wrong than correct responses, they generate fewer consistently wrong responses compared to consistently correct ones.
Figure 6: Comparison between llama3instruct-8b and llama2chat-7b (zero-shot setting). The former model is slightly less factual according to NCR, but more consistent in its responses (see NCCR, $C_{correct}$, and $C^{s}_{wrong}$). We observe similar trends in the four-shot and with two unsure shots settings.
Comparisons Zero-Shot Four-Shot Four-Two
NCR vs. UR 0.27 0.34 0.62
NCCR vs. IUR --0.17 --0.12 0.41
$C_{\textit{correct}}$ vs. $C_{\textit{wrong}}$  0.81  0.78 0.51
Table 2: Pairwise correlation of LLM performance on different metrics under different prompt settings. The correlation values are computed across all LLMs (26 data points). We report Pearson’s $\rho$, diacritic  denotes statistical significance ($p<0.05$). Four-Two refers to the four-shot setting with two unsure shots.
Performance on seen knowledge is not predictive of performance on unseen knowledge.
In Table 2, we examine whether metrics applied to seen knowledge can be used to extrapolate model performance on unseen knowledge by reporting correlation values (Pearson’s $\rho$) between NCR and UR, and NCCR and IUR. As can be seen, in both zero- and four-shot settings, correlations are not statistically significant. Nevertheless, correlations are significant in all metric comparisons in the four-shot with two unsure shots setting. While in general model performance with seen knowledge does not transfer to unseen knowledge, specific prompt manipulations can enhance the correlation between metrics (see last column in Table 2).
LLMs are consistently right and wrong!
Table 2 reports a positive, significant correlation between $C_{correct}$ and $C_{wrong}$. This result implies that LLMs demonstrating high consistency in correct responses also tend to exhibit high consistency in wrong responses. This finding contradicts our expectation of high $C_{\textit{correct}}$ and low $C_{\textit{wrong}}$ and highlights a notable flaw in current models. Future work should address this issue, e.g., by instructing LLMs to achieve the desired consistency behavior.
Figure 7: The impact of model size on LLM performance, measured with NCR, UR, Ccorrect, Cwrong, NCCR, and IUR. Different metrics are color-coded. LLMs are shown in three sizes, small, medium, and large and are grouped into ‘base’ and fine-tuned ones.
5 Analysis
In this section, we explore the effects of model size, fine-tuning, and ICL on LLM performance based on different metrics. We also analyze how LLMs handle unseen knowledge by looking at the distribution of uninformative responses and the impact of different question types.
Figure 8: The impact of fine-tuning on LLM performance, measured with NCR, UR, Ccorrect, Cwrong, NCCR, and IUR. Different metrics color-coded. This analysis only considers the performance of Llama2, Llama3, Mistral, and Gemma as these families include both base LLMs and fine-tuned versions. Models are shown in three sizes, small, medium, and large.
Figure 9: The impact of ICL on LLM performance, measured with NCR, UR, Ccorrect, Cwrong, NCCR, and IUR. Different metrics are color-coded. We compare zero-shot and four-shot settings; and zero-shot against four-shot with two unsure shots. LLMs (‘base’ and fine-tuned ones) are in three sizes, small, medium, and large.
Figure 10: The impact of question type on LLM performance as measured by Uninformative Rate (UR) on unseen knowledge. Questions are grouped into types based on answer they require. Higher values have darker shades.
5.1 The Impact of Model Size
Larger LLMs perfom better on seen knowledge but worse on unseen knowledge.
Figure 7 demonstrates that as model size increases, both NCR (blue line) and NCCR (purple line) improve. This trend indicates that larger LLMs perform better on questions about seen knowledge. However, for base LLMs in both zero-shot and four-shot settings, IUR (brown line) decreases as model size increases. This trend suggests that as LLMs become larger, they become more consistent at delivering wrong responses for questions regarding unseen knowledge.
Larger LLMs are more consistent, even with wrong responses.
In Figure 7, we observe that as model size increases, both $C_{\textit{correct}}$ (green line) and $C_{\textit{wrong}}$ (red line) increase significantly. While higher consistency scores for correct responses are expected and useful, the higher consistency scores for wrong responses pose a potential risk. Larger models may consistently produce convincing but wrong information, which could lead to misinformation if not carefully managed.
5.2 The Impact of Fine-tuning
Fine-tuning improves performance on unseen knowledge but negatively affects performance on seen knowledge.
As illustrated in Figure 8, UR (orange line) and IUR (brown line) show significant improvement in LLMs after fine-tuning, indicating an enhanced ability to distinguish and respond to unseen knowledge appropriately. However, the decreasing NCCR (purple line) suggests they become worse at handling seen knowledge after fine-tuning.
Fine-tuning does not make LLMs more consistent.
As depicted in Figure 8, there is no noticeable increase in $C_{\textit{correct}}$ (green line) after fine-tuning. In fact, fine-tuning even has a negative impact on $C_{\textit{correct}}$ for medium-sized LLMs. These results indicate that current fine-tuning techniques fail to enhance the consistency of correct responses. Furthermore, $C_{\textit{wrong}}$ (red line) does not decrease after fine-tuning either, which suggests that fine-tuning also fails to reduce the model’s presistence on wrong responses.
5.3 The Impact of ICL
Unsure shots improve LLM performance on unseen knowledge.
As shown in Figure 9, incporating two unsure shots in the four-shot setting substantially increases UR (orange line) and IUR (brown line) across model sizes. In contrast, performance on unseen knowledge deteriorates when using four-shots only for all sizes of base LLMs and small fine-tuned LLMs, compared to their zero-shot counterparts.
ICL does not improve LLM perfomance on seen knowledge.
Figure 9 demonstrates that ICL does not improve LLM performance according to NCR (blue line) or NCCR (purple line) metrics. For small/medium base LLMs and small fine-tuned LLMs in the four-shot setting, ICL even decreases NCR performance. This finding contrasts with previous results about the ability of LLMs to learn from a few examples Brown et al. (2020); Chada and Natarajan (2021); Touvron et al. (2023); Bai et al. (2023). Earlier work has mostly focused on the correct rate of LLMs in few-shot settings without unsure shots. Based on the results presented in Table 9 in Appendix D, we observe that for most LLMs, the four-shot setting indeed substantially improves the correct rate but also increases the wrong rate. Incorporating unsure shots helps decrease the wrong rate but also decreases the correct rate.
ICL makes LLMs less consistent, even on correct responses!
Figure 9 shows that ICL leads to reductions in both $C_{\textit{correct}}$ (green line) and $C_{\textit{wrong}}$ (red line) for medium/large fine-tuned LLMs and large base LLMs. While a lower $C_{\textit{wrong}}$ is desirable, the decrease in $C_{\textit{correct}}$ is concerning. This indicates that ICL negatively affects the reliability of LLMs at providing correct responses.
5.4 Model Behavior on Unseen Knowledge
Base LLMs overestimate their knowledge on numerical and temporal questions.
Figure 10 reports UR performance across models, broken down per question type (e.g., number, person, location). As can be seen, most base LLMs exhibit significantly lower UR performance on questions requiring numerical or temporal responses. This indicates that base LLMs often provide misleading information for such questions, even when they lack the relevant knowledge.
Fine-tuned LLMs can explicitly admit when they don’t know.
As illustrated in Figure 11 (fourth column), fine-tuned LLMs are able to explicitly acknowledge their lack of knowledge by answering ‘unsure’ (see definition in Section 3.3) to questions about unseen knowledge. In contrast, base LLMs often produce responses classified as ‘none’ or ‘repetition’ in the absence of unsure shots (contrast columns one and two with column three in Figure 11).
6 Related Work
Petroni et al. (2019) were the first to introduce the concept of using pre-trained language models as knowledge bases. They introduced LAMA, a benchmark with questions structured as "fill-in-the-blank" cloze statements, and found that BERT Devlin et al. (2019) retains relational knowledge competitive with traditional NLP methods that have some access to oracle knowledge.
Roberts et al. (2020) were the first to measure the extent to which language models trained on unstructured text can implicitly store and retrieve knowledge using natural language queries. Specifically, they fine-tuned T5 Raffel et al. (2019) to answer questions without access to any external context or knowledge and showed that this approach performs competitively with open-domain systems that explicitly retrieve answers from an external knowledge source when answering questions. On a similar vein, Wang et al. (2021) fine-tuned BART Lewis et al. (2020) with related passages to instill factual knowledge and used the accuracy of masked span recovery to measure the extent to which this knowledge was memorized by the model. They found it was challenging for pre-trained language models like BART to remember facts seen in training, and answer questions, even in cases where the relevant knowledge was retained. He et al. (2024) explicitly trained T5 Raffel et al. (2019) and LLAMA2 Touvron et al. (2023) to memorize world knowledge from Wikidata on a large scale and then used exact match and F1 scores to evaluate knowledge retention. Their results showed LLMs hold promise as large-scale KBs capable of retrieving facts and responding with flexibility, but are less proficient at inferring new knowledge through reasoning.
Sun et al. (2023) proposed a benchmark consisting of 18,000 question-answer pairs representing facts with varying popularity (i.e., high, medium, low) and assessed the knowledge retained by various language models on this benchmark. Unlike previous studies, Sun et al. (2023) reported both accuracy (the percentage of questions answered correctly) and hallucination rate (the percentage of questions answered wrongly). They showed models are particularly bad at answering questions from medium- and low-frequency facts.
Our work also examines whether LLMs can be used in lieu of more traditional KBs. Compared to previous research, we proposed a comprehensive evaluation framework which assesses not only the ability of LLMs to recall seen knowledge but also their ability to respond in the face of unseen knowledge. In addition, we evaluate LLM consistency when answering questions about identical knowledge.
7 Conclusion
In this paper, we defined a set of criteria that LLMs functioning as KBs should meet, focusing on factuality and consistency. We proposed various metrics operationalizing these criteria and used them to assess LLM performance when answering questions pertaining to both seen and unseen knowledge. We evaluated 26 popular LLMs and found that gpt-3.5-turbo is the most reliable among them. Additionally, we examined the impact of model size, fine-tuning, and ICL. Fine-tuning and ICL with unsure shots were shown to improve LLM capabilites when providing responses to questions about unseen knowledge, but do not significantly improve performance on seen knowledge. Increasing model size boosts performance on seen knowledge but at the expense of performance on unseen knowledge. Notably, neither ICL nor fine-tuning were successful at improving LLM consistency. This highlights the critical need for continued research to develop more robust strategies that ensure both factuality and consistency, enabling LLMs to effectively function as KBs.
Appendix A Datasets
A.1 SeenQA Dataset
SeenQA is composed of questions selected from the following datasets:
  1. 1.
  2. 2.
  3. 3.
Wikipedia is a common source in the pre-training data of large language models (LLMs). Comparing the knowledge cutoff dates provided in Table 4, we can deduce that the knowledge involved in these three datasets must have been seen during training by the LLMs used in our study. SeenQA contains 3000 questions selected through a three-step process:
  1. 1.
    Factoid Question Extraction: We filter out questions starting with "why," those with multiple different answers, or answers longer than five tokens.
  2. 2.
    Removal of Time-Sensitive Questions: We use gpt-4-1106-preview to detect questions whose answers may change over time, using the prompt shown in Table 5. Such questions are then removed from the dataset.
  3. 3.
    Sampling: Finally, we randomly sample 1,000 questions from each dataset’s test or development set, if the test set is unavailable, discarding the supporting context to adapt to a closed-book setting.
A.2 UnseenQA Dataset
UnseenQA was created using 20 templates, shown in Table 3, covering five answer types: number, person, time, location, and others. These templates wet=re hand-written by the authors. Templates T1, T2, T5, T6, and T7 are about future events, so their answers are unknown at the time of writing. The remaining templates involve fictional persons whose names and related information cannot be found on the Internet. The templates include three types of placeholders:
  1. 1.
    Country/Region: We collected 150 country/region names from the list of National Olympic Committees on a Wikipedia page.
  2. 2.
    Medal Event: We collected 150 medal events from the official programme of the Olympic games, Paris 2024.
  3. 3.
    Person: We randomly sampled 150 person names from combinations of 100 first names, 100 middle names, and 100 last names. We manually checked these names on Google to ensure they have no Internet presence.
UnseenQA consists of 3000 questions, with 150 questions per template. This dataset was created on April 13, 2024. Therefore, LLMs trained on knowledge before this date should not know the answers.
@@ A.2 UnseenQA Dataset
Appendix B LLMs used in this project

### v2

@@ 3.1 LLM Selection
We evaluate 26 popular LLMs, including gpt-3.5-turbo, Flan-t5, llama1, llama2, llama3, mistral, gemma, and Phi2. Detailed descriptions of the evaluated LLMs are provided in Table 4 in Appendix B. Detailed descriptions are available in Table 4 in Appendix B. We test LLMs of different sizes: small (0.08B–3B), medium (7B–13B), and large (65B–70B). "Fine-tuned LLMs" refer to those fine-tuned via instruction-tuning or reinforcement learning from human feedback (e.g., llama3instruct-8B), while "base LLMs" refer to models without fine-tuning (e.g., llama3-8B).
@@
3.2 Datasets
SeenQA
SeenQA is a composite dataset comprising 3,000 questions sourced from the test sets (or development sets, where test sets were unavailable) of Natural Questions, TriviaQA, and PopQA (see Appendix A). All three datasets are derived from Wikipedia, with a knowledge cutoff date no later than December 2018. Given that Wikipedia is a frequent source for pre-training LLMs and the evaluated LLMs in this study have a knowledge cutoff date beyond April 2019 (as shown in Table 4), it can be inferred that the knowledge in these datasets is available to LLMs during their training. The creation of SeenQA involved a three-step process: 1) Factoid Question Extraction: exclude "why" questions, those with multiple answers, or answers exceeding five tokens. 2) Time-Sensitive Question Removal: use gpt-4-1106-preview (prompt in Table 5) to detect and remove questions with time-variant answers. 3) Random Sampling: randomly select 1,000 questions per source, discarding supporting context for a closed-book setting.
UnseenQA
UnseenQA is a new QA dataset designed to ensure the LLMs tested in this study lack access to knowledge required to answer questions. It consists of 3,000 questions generated from 20 hand-written templates (150 questions per template), as detailed in Table 1, spanning five answer types: number, person, time, location, and others. Templates T1–T7 focus on future events with unknown answers at the time of writing, while the remaining templates involve fictional individuals whose names and details are not available online. The templates feature three placeholder types: 1) Country/Region: 150 names sourced from the National Olympic Committees listed on the Wikipedia page. 2) Medal Event: 150 medal events from the official programme of the Olympic games, Paris 2024. 3) Person: 150 randomly generated names from combinations of 100 first, middle, and last names, manually verified to lack online presence. The dataset was created on April 13, 2024. All LLMs studied were released before this date and thus lacked access to the knowledge.
3.3 Evaluation on a Single Response
Uninformative
We classify uninformative responses into three categories: repetition, none, and unsure (see Appendix C). Repetition refers to responses that repeatedly echo a specific string. We detect this using regular expressions and word frequency analysis. None includes responses that lack relevant content, such as empty strings or those merely repeating the question. Unsure applies to responses where the model says it is unable to answer or does not know. Examples include phrases like "I am not sure" or "I am just an AI", etc.
Correct
A response is considered correct if it is an exact match with the ground truth or is similar enough as judged by GPT-4o. Following prior work Sun et al. (2023), which reported a 98% agreement rate between ChatGPT and human evaluations, we adopt their evaluation prompt (detailed in Table 7 in Appendix D).
Figure 2: Top 15 LLMs ranked by (a) Net Correct Rate, (b) Uninformative Rate, (c) $C_{\textit{correct}}$, (d) $C_{\textit{wrong}}$, (d) Net Consistent Correct Rate, (e) Inconsistent/Uninformative Rate. Values are scaled by 100 (full results in Appendix E).
Consistency Score
To compute $\text{Cons}(q,r)$, we set $X_{\text{MCQs}}$ (total number of MCQs) to 20, and each MCQ includes question $q$ and 5 options (the original response $r$, 3 random distractor options, and an ‘unsure’ option).
3.4 Prompts and Hyper-parameters
All LLMs were evaluated using greedy decoding (temperature 0 for gpt-3.5-turbo) with a maximum of 100 new tokens. To provide a comprehensive evaluation, we experimented with three types of prompt settings: zero-shot, four-shot, and four-shot with two unsure shots. To avoid any bias introduced by fixed examples, we employed a dynamic few-shot method following the work of Nori et al. (2023). For further details of the prompt settings, refer to Appendix D.
4 Quantitative Results
We present results for all LLMs under different prompt settings in in Appendix E (Table 9: factuality, Table 10: consistency, and Table 11: reliability). LLM rankings based on different metrics are shown in Figure 6 and Figure 7, and also in Appendix E. In the remainder, we discuss our results in dynamic four-shot settings.
Which LLM is most factual?
As shown in Figures 2 (a) and (b), gpt-3.5-turbo performs best on seen knowledge, while gemma-instruct-7B excels at unseen knowledge. gpt-3.5-turbo and llama2chat-70b emerge as the most balanced models, demonstrating top performance across both seen and unseen knowledge. Notably, Figure 2 (a) highlights a limitation of CR, the standard metric for factuality. For instance, while llama1-13b has a significantly higher CR than gemma-instruct-7B, their behavior differs in terms of wrong responses. llama1-13b exhibits a 33.50% higher WR rate, leading to a lower NCR compared to gemma-instruct-7B.
Which LLM is most consistent?
As shown in Figures 2 (c) and (d), llama3instruct-70b is most consistent on questions it answers correctly, while llama2chat-7b is least consistent on questions it answers incorrectly. Among these models, llama3instruct-8B is most balanced, with a high $C_{\textit{correct}}$ and low $C_{\textit{wrong}}$.
Which LLM is most reliable?
As shown in Figures 2 (e) and (f), gpt-3.5-turbo excels at seen knowledge, while gemma-instruct-7B performs best on unseen knowledge. Among these models, gpt-3.5-turbo demonstrates balanced performance, with a top NCCR and a high IUR. Note that a more factual LLM is not necessarily more reliable. For example, llama3-instruct-8B ranks below the 15th in NCR and 13th in NCCR due to its relatively high $C_{\textit{correct}}$ and low $C_{\textit{wrong}}$.
5 Discussion
5.1 Correlation Analysis
Comparisons Zero-Shot Four-Shot Four-Two
NCR vs. UR 0.27 0.34 0.62
NCCR vs. IUR --0.17 --0.12 0.41
$C_{\textit{correct}}$ vs. $C_{\textit{wrong}}$ 0.81 0.78 0.51
NCR vs. $C_{\textit{correct}}$ 0.60 0.41 0.37
NCR vs. $C_{\textit{wrong}}$ 0.48 0.64 0.41
UR vs. $C_{\textit{correct}}$ 0.43 0.07 0.35
UR vs. $C_{\textit{wrong}}$ 0.34 0.05 0.48
Table 2: Pairwise correlation of LLM performance on different metrics under different prompt settings. The correlations are computed across all LLMs (26 data points). We report Pearson’s $\rho$, with underlined values indicating statistical significance ($p<0.05$). Four-Two refers to the four-shot setting with two unsure shots.
Does an LLM with strong performance on seen knowledge tend to perform strong on unseen knowledge?
We next examine whether there is a linear relationship between performance on seen and unseen knowledge. Performance on seen knowledge does not reliably predict performance on unseen knowledge in zero-shot and few-shot settings (without unsure shots). Table 2 (the first and second rows) shows the correlation values (Pearson’s $\rho$) between NCR and UR, and NCCR and IUR, revealing no statistically significant correlations in these settings. However, in the four-shot setting with two unsure shots, correlations across all metrics are significant. While seen knowledge performance does not transfer to unseen knowledge, specific prompt manipulations can improve these correlations (see the last column in Table 2).
Does an LLM with high consistency score on correctly answered questions tend to perform less consistent on wrongly answered questions?
As shown in Figure 11 and Table 10 in Appendix E, an LLM tends to be more consistent on questions it answers correctly than on those it answers wrongly. However, when comparing different LLMs, models with higher $C_{\textit{correct}}$ also tend to have higher $C_{\textit{wrong}}$, as indicated by the positive, significant correlation reported in Table 2 (the third row). This finding suggests that LLMs consistent with correct responses are also consistent with wrong responses, contradicting the expectation of high $C_{\textit{correct}}$ and low $C_{\textit{wrong}}$. This highlights a notable flaw in current models, which future work should address.
Does strong factuality performance correlate with strong consistency performance?
As shown in Table 2 (last four rows), there is a positive correlation between NCR/UR and $C_{\textit{correct}}$/$C_{\textit{wrong}}$, particularly in zero-shot settings. This suggests that LLMs with strong factuality performance tend to be confident in their responses, maintaining consistency whether those responses are correct or wrong.
5.2 Model Size, Fine-tuning and ICL Impact
Figure 3: The impact of (a) model size, (b) fine-tuning, (c) ICL on LLM performance, measured with NCR, UR, Ccorrect, Cwrong, NCCR, and IUR. Different metrics are color-coded. See Appendix E for more detailed visualization. Values are scaled by 100.
Figure 4: The impact of question type on LLM performance, measured by Uninformative Rate (UR) on unseen knowledge (scaled by 100). Questions are grouped by answer type. Higher values have darker shades.
How does model size affect LLMs performance?
As shown in Figure 3(a), larger LLMs perform better on seen knowledge but worse on unseen knowledge. As model size increases, both NCR (blue line) and NCCR (purple line) improve, indicating better performance on questions related to seen knowledge. However, while the UR (orange line) increases, the IUR (brown line) decreases. This suggests that larger models make fewer wrong responses on unseen knowledge but tend to give more consistent wrong responses. We also observe that larger LLMs are more consistent, even with wrong responses. Both $C_{\textit{correct}}$ (green line) and $C_{\textit{wrong}}$ (red line) rise significantly with model size. While higher consistency in correct responses is desirable, the increase in consistency for wrong responses presents a risk. Larger models may confidently and consistently produce incorrect yet convincing information, heightening the potential for misinformation if not carefully managed.
How does fine-tuning affect LLMs performance?
Figure 3(b) suggests that fine-tuning improves performance on unseen knowledge but degrades performance on seen knowledge. After fine-tuning, both UR (orange line) and IUR (brown line) increase significantly, indicating improved handling of unseen knowledge. However, the decline in NCR (blue line) and NCCR (purple line) shows that fine-tuning reduces the model’s effectiveness with seen knowledge. We also see that fine-tuning has no impact on consistency. There is no notable change in $C_{\textit{wrong}}$ (red line) or $C_{\textit{correct}}$ (green line) after fine-tuning which suggests that current instruction-tuning and RLHF methods do not improve LLM consistency.
How does ICL affect LLMs performance?
Figure 3(c) shows that ICL does not improve performance on seen knowledge, but unsure shots enhance performance on unseen knowledge. ICL does not increase NCR (blue line) or NCCR (purple line), which contrasts with previous findings on LLMs’ ICL ability Brown et al. (2020); Chada and Natarajan (2021); Touvron et al. (2023); Bai et al. (2023). Prior research largely focuses on the correct rate in few-shot settings without unsure shots. As shown in Table 9 and Figure 12 (Appendix E), while the four-shot setting significantly increases the correct rate, it also raises the wrong rate. Adding unsure shots helps reduce the wrong rate but also lowers the correct rate. For unseen knowledge, incorporating two unsure shots in the four-shot setting substantially improves UR (orange line) and IUR (brown line), indicating better handling of unknown questions. ICL with unsure shots reduces consistency in wrong responses. Adding two unsure shots in the four-shot setting decreases $C_{\textit{wrong}}$ (red line). Without unsure shots, there is no significant change in $C_{\textit{wrong}}$ (red line) or $C_{\textit{correct}}$ (green line).
5.3 Behavior on Unseen Knowledge
How do LLMs perform on different types of questions about unseen knowledge?
Figure 4 shows that base LLMs tend to overestimate their knowledge when answering numerical and temporal questions. Their UR scores are significantly lower for these types of queries, indicating a tendency to provide misleading responses even when they lack relevant knowledge.
6 Related Work
Petroni et al. (2019) first explored using pre-trained LMs as KBs, introducing LAMA and showing that BERT retains relational knowledge with precision as the metric. Roberts et al. (2020) evaluated knowledge storage and retrieval using natural language queries, measuring accuracy. Wang et al. (2021) fine-tuned BART with related passages to instill factual knowledge, assessing masked span recovery accuracy. He et al. (2024) trained T5 and LLaMA2 on Wikidata to evaluate large-scale knowledge retention via exact match and F1 scores. Sun et al. (2023) tested LLMs on 18,000 fact-based QA pairs, reporting both accuracy and hallucination rates.
Many studies have highlighted the issue of factuality in current LLMs. To address this, new benchmarks have been proposed to assess factuality Muhlgay et al. (2024); Zhao et al. (2024); Liu et al. (2024), although these still rely on correct rate as the primary metric. Several methods to enhance LLM factuality have also been introduced Wang et al. (2022); Hase et al. (2024); Cohen et al. (2024); Qin et al. (2024). However, evaluating these improved models falls outside the scope of our experiments and is suggested for future work.
Research on consistency Rajan et al. (2024); Sreekar et al. (2024); Saxena et al. (2024) has shown that LLMs often struggle with providing consistent responses. These works, however, do not differentiate between the consistency expectations for correctly and wrongly answered questions.
Compared to previous research, we propose a comprehensive framework to evaluate not only whether LLMs recall seen knowledge but also their ability to respond to unseen knowledge. In addition, we evaluate LLM consistency when answering questions about identical knowledge.
7 Conclusion
In this paper, we rethink the requirements for evaluating LLMs as KBs and propose criteria emphasizing factuality and consistency, and the combination thereof which we argue is an indicator of reliability. We proposed various metrics operationalizing these criteria and used them to assess LLM performance when answering questions pertaining to both seen and unseen knowledge. We evaluated 26 LLMs on our newly proposed SeenQA and UnseenQA datasets, and examined the impact of model size, fine-tuning, and ICL. Our experimental results highlight the critical need for continued research to develop more robust strategies that ensure both factuality and consistency, enabling LLMs to reliably function as KBs.
Limitations
First, due to budget constraints, we conducted in-depth evaluations on only a single closed-source LLM (GPT-3.5-turbo). Nevertheless, this limitation does not diminish the contributions of our work. Our study includes a broad comparative analysis of 26 different LLMs, ensuring that the insights gained are comprehensive and not confined to any single model. Furthermore, the primary objective of our paper is to introduce a systematic framework for evaluating LLM-as-KB. This framework is universally applicable, offering value for the evaluation of a wide range of LLMs beyond those specifically analyzed in this study.
Second, our evaluation focuses primarily on factoid questions, which assess the models’ ability to recall specific factual knowledge rather than perform complex reasoning. Investigating how LLMs handle such complex queries remains an important direction for future research.
Appendix A Datasets
SeenQA is composed of questions selected from the following open-sourced datasets:
  1. 1.
  2. 2.
  3. 3.
Wikipedia is a common source in the pre-training data of large language models (LLMs). Comparing the knowledge cutoff dates provided in Table 4, we can deduce that the knowledge involved in these three datasets must have been seen during training by the LLMs used in our study.
Uninformative Type Responses Examples
repetition
1. (a) (b) (c) (d) (e) (f) (g) (h) (i) (j) (k) (l) (m) (n) (o) (p) (q) (r) (s) (t) (u) (v) (w) (x) (y) (z)
(aa) (ab) (ac) (ad) (ae) (af) (
1. 2. 3. 4. 5. 6. 7. 8. 9. 10. 11. 12. 13. 14. 15. 16. 17. 18. 19. 20. 21. 22. 23. 24. 25. 26. 27.
a swollen swollen swollen swollen swollen swollen swollen swollen swollen swollen
swollen swollen swollen swollen swollen swollen swollen swollen
none
noah catherine cooper’s mom is ________.
<a href="https://i.stack.imgur.com/88888.png" rel="nofollow noreferrer"><image></a>
unsure
unsure
i do not know.
i’m just an ai, i don’t have access to real-time information or the ability to predict the future
Table 3: Uninformative Responses Examples
Appendix B LLMs Used